OpenStudy (anonymous):

At a temperature of 320K, the gas in a cylinder has a volume of 40.0 liters. If the volume of the gas is decreased to 20.0 liters, what must the temperature be for the gas pressure to remain constant? A. 160 K B. 273 K C. 560 K D. 140 K

OpenStudy (cwrw238):

V1/T1 = V2/T2 40/320 = 20/T2 solve for T2

OpenStudy (anonymous):

560?

OpenStudy (cwrw238):

cross multiplying 40T2 = 20*320 T2 = (20*320) / 40 = 160

OpenStudy (anonymous):

I got it, thank you!